Published in the Proceedings of the , pages –

A New Intuitionistic Proof of Usability of the Recommended Style File for the ACM

Conference Papers

Jürgen Vollmer^{∗}

Karlsruhe

Juergen.Vollmer@acm.org Mickey Mouse

Enthausen University

Mickey.Mouse@entenhausen.org

ABSTRACT

This document demonstrates how to use the L^{A}TEX2e acmconf class by exhibiting itself as an example. You are expected to be familiar with [1]. The best way to use this file is to use it as a template, i.e., replace the prose in it by your own^{1}.

KEYWORDS

L^{A}TEX2e-class, ACM proceedings

1 INTRODUCTION

To understand this file read the source and not the typeset version. If you are reading this in the typeset version you might as well stop — it is not supposed to make sense.

2 THE STORY BEGINS…

A real article is supposed to have some deep results and good explanations. That, however, is your job and not mine so you should replace this text with something more appropriate^{2}..

∗Happy L^{A}TEXing 1And may use footnotes. 2Another a footnote

3 SOME OFTEN USED L^{A}TEX COM-MANDS

3.1 emph, etc.

Text may be set as emph. Text may be set as texttt. Text may be set as \underline{unterline}. Text may be set as textbf. Text may be set as textrm. Text may be set as tiny. Text may be set as scriptsize. Text may be set as footnotesize. Text may be set as normalsize. Text may be set as large. Text may be set as Large. Text may be set as LARGE. Text may be set as huge.

Text may be set as Huge. Text may havesuper andsubscripts.

3.2 itemize

• More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references.

• More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references.

3.3 enumerate

1. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references.

2. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references.

Nice Postscript, isn’t it?

Same, a little bit smaller:

Figure 1: This is a nice floating figure

3.4 description

Foo More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references.

Bar More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references.

3.5 center and tabular

left  center  right
AAAAAAAA BBBBBBBB CCCCCCCC
AAAAAAAA BBBBBBBB CCCCCCCC
AAAAAAAA BBBBBBBB CCCCCCCC
AAAAAAAA BBBBBBBB CCCCCCCC
AAAAAAAA BBBBBBBB CCCCCCCC
AAAAAAAA BBBBBBBB CCCCCCCC
1  2

3.6 figure and Postscript pictures

Have a look to to figure 1 and 2.

4 THE STORY CONTINUES 1

This is a \section.

More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references.

More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references.

4.1 The Story Continues 2

This is a \subsection.

More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references.

More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references.

4.1.1 The Story Continues 3

This is a \subsubsection.

More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references.

More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references.

4.1.1.1 The Story Continues 4

This is a \subsubsubsection.

More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references.

More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text This figure uses both columns, using figure*

Figure 2: This is a nice floating figure

should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references.

4.1.1.1.1 The Story Continues 5

This is a \subsubsubsubsection.

More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references.

More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references.

The Story Continues 6 This is a \paragraph. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references.

More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references.

The Story Continues 7 This is a \subparagraph. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references.

More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references.

More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references. More text should follow, but keep in mind that a limit of 6 pages has been set, including figures and references.

5 CONCLUSION

The end, at last! In this example there really are no results or points to summarize but I trust your article has more food for though and thus will need a conclusion.

A APPENDICES

If you have any, appendices might go here. Note that appendices should not be used to circumvent the word count limit.

This is ”doing it by hand” — you might be better off using BibTeX.

REFERENCES

[1] Leslie Lamport: L^{A}TEX, A Document Preparation System, Addison Wesley 1994.